OpenStudy (anonymous):

A six-sided number cube with faces numbered 1, 2, 3, 4, 5, and 6 is rolled. What is the probability of rolling a 6 or a number less than 3? A. 1/18 B. 1/6 C. 1/3 D. 1/2

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

Should be D

OpenStudy (anonymous):

how did you get d

OpenStudy (anonymous):

3/6=1/2

OpenStudy (anonymous):

Im working on this in math

OpenStudy (anonymous):

A six-sided number cube is rolled. Find the probability of rolling a 3 or a number greater than 4. A. B. C. D.

OpenStudy (anonymous):

@Anthonyman1011123

OpenStudy (anonymous):

3/6=1/2

OpenStudy (anonymous):

a.1/6 b.1/2 c.2/3 d.1/18

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Anything else?

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

A box contains twenty cards numbered 1 through 20. One card is to be randomly drawn from the box. What is the probability that the card drawn will show a multiple of 6, a multiple of 7, or a multiple of 8? A.1/2 B.7/20 C.1/3 D.1/20

OpenStudy (anonymous):

7/20 i believe. 3 cards are multiples of 6 (6,12,and18) 2 cards are multiples of 7 (7and14) and 2 cards are multiples of 8 (8and16)

OpenStudy (anonymous):

so 7/20

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i gave you a medal @Anthonyman1011123

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

and i fanned you